OpenStudy (anonymous):

Given f(x)=10x+8 and g(x)=1x+4 then what is f(g(3)) ?

jimthompson5910 (jim_thompson5910):

first we need the value of `g(3)`

jimthompson5910 (jim_thompson5910):

g(x) = 1x+4 g(3) = 1*3+4 ... replace every x with 3 g(3) = ______ (fill in the blank)

OpenStudy (anonymous):

f(x)=10x+8 and g(x)=1x+4

jimthompson5910 (jim_thompson5910):

yes you already posted that

OpenStudy (anonymous):

id understand

jimthompson5910 (jim_thompson5910):

g(x) = 1x+4 g(3) = 1*3+4 ... replace every x with 3 g(3) = ______ (fill in the blank, tell me what you get here)

OpenStudy (anonymous):

7

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so we can replace the `g(3)` with `7` since they are equivalent

jimthompson5910 (jim_thompson5910):

`f(g(3)) ` will turn into `f(7)`

jimthompson5910 (jim_thompson5910):

to compute f(7), you follow similar steps done to compute g(3)

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

38

jimthompson5910 (jim_thompson5910):

38 is not correct

jimthompson5910 (jim_thompson5910):

you might have plugged in x = 3 by mistake

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

go back up and see why we're plugging in x = 7 instead

OpenStudy (anonymous):

is it because g(x) was 7

jimthompson5910 (jim_thompson5910):

it's because the output g(3) is equal to 7

OpenStudy (anonymous):

alright

jimthompson5910 (jim_thompson5910):

anyways, what is the value of f(7) ?

OpenStudy (anonymous):

78

jimthompson5910 (jim_thompson5910):

correct, that is the final answer

OpenStudy (anonymous):

thanks